Simplify.

square root of 72

square root of 50x to the 6th power y to the 5th power

Found 2 solutions by stanbon, checkley71:
Answer by stanbon(48516)

Simplify.
square root of 72 = sqrt(36*2) = 6sqrt2
---------------------
square root of 50x to the 6th power y to the 5th power
sqrt(50x^6y^5) = sqrt[(25x^6y^4)*(2y)] = (5x^3y^2)*sqrt(2y)
============
Cheers,
Stan H.

Answer by checkley71(8403)

SQRT72=SQRT(36^2)=6SQRT2 ANSWER.
----------------------------------
SQRT50X^6Y^5=SQRT(25*2)(X^6Y^5)=(5X^3Y^2)SQRT2Y ANSWER.